Seminar on derived categories in algebraic geometry

Thursday, February 16, 3.00-4.00, MIT, room 2-151.

Braid group actions arising from noncommutative resolutions.

by Roman Bezrukavnikov (MIT).

I will briefly describe the affine braid group actions (and possible generalizations) on derived categories of coherent sheaves on some algebraic

symplectic varieties arising from semisimple Lie algebras (this has been discussed also, from a different perspective, in my talk in Kac’s seminar in

the fall). In this talk I will try to point out the relation to works of Bridgeland and Seidel-Thomas discussed in this seminar in the fall (though

the talk will not rely on this material).

The seminar is a continuation of the seminar from the fall, though new participants are most welcome, the material discussed in the fall is not a

prerequisite.

NB: We start with the first talk on Thursday, 2/16, but starting next week will most probably meet on Tuesdays, 5-6:30 pm.

seminar contacts: Roman Bezrukavnikov (bezrukav@math.mit.edu); Rina Anno (rina@math.harvard.edu)